OpenStudy (anonymous):

Find the line that passes through the point (0, 2) and is parallel to the line y = -7/3x - 4.

OpenStudy (nali):

lines are parallel when they have the same slope so use the slope that u have -7/3 then plug in 0and 2 as x and and y respectively in this equation Y-y= m (X-x) Y-2= -7/3 (X-0) Y-2 = -7/3X move the 2 to he other side and u will get ur equaiotn

OpenStudy (anonymous):

so its y=-5/3x

OpenStudy (nali):

no u can not add the 2 to the -7/3 because it has an x next to it but it u just have to move it to ther side like this y=-7/3x+ 2

OpenStudy (anonymous):

these are the answers y = -7/3x + 2 y = 7/3x - 4 y = 3/7x - 4

OpenStudy (anonymous):

oh duh

OpenStudy (anonymous):

thanks alot

OpenStudy (nali):

ur welcome

OpenStudy (anonymous):

thxs